OpenStudy (phagey):

Calculate the time required for an object to increase velocity from 25 ft/s to 125 ft/s. Use the formula V = V0 + 32t, where V0 is the starting velocity of a falling object, V is the final velocity of the falling object, and t is the time for the object to fall. 3.125 seconds 4800 seconds 4.69 seconds 3200 seconds

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle v=v_0+32t }\)

OpenStudy (solomonzelman):

All you need to do is to plug in your numbers. The final velocity is 125 ft/s, and the initial velocity is 25 ft/s. \(\color{#000000 }{ \displaystyle v=125~{\rm ft/s} }\) \(\color{#000000 }{ \displaystyle v_0 =25~{\rm ft/s} }\)

OpenStudy (solomonzelman):

Then, just solve for time (t).

OpenStudy (phagey):

ok thank you